# Elevated UV luminosity density at Cosmic Dawn explained by non-evolving, weakly-mass dependent star formation efficiency

### v1

@@ 3.4 A theoretical model for the UV luminosity function and density at $z>6$
In the left panel of Fig. 10, we plot ${\rm d}\rho_{\rm UV}/{\rm d}\lg{}M_{\rm halo}$ as function of mass and redshift as predicted by our theoretical model for a mass-dependent scatter. According the figure, the majority of the cosmic UV luminosity at $z\sim{}6-15$ is produced by galaxies in halos of moderate mass ($M_{\rm halo}\sim{}10^{9}-10^{11}\,M_{\odot}$). For instance, at $z=6$ half the UV luminosity arises in halos with $M_{\rm halo}\sim{}10^{10.3}-10^{11}\,M_{\odot}$ if a UV magnitude limit of $M_{\rm UV,lim}=-17$ is chosen. The mass range of halos responsible for the majority of the UV luminosity shifts towards lower masses with increasing redshift, e.g., it is $M_{\rm halo}\sim{}10^{9.1}-10^{9.6}\,M_{\odot}$ at $z=15$. For $M_{\rm UV,lim}=-14$, the corresponding halo mass ranges are $M_{\rm halo}\sim{}10^{9.7}-10^{10.7}\,M_{\odot}$ and $M_{\rm halo}\sim{}10^{8.8}-10^{9.3}\,M_{\odot}$. The downward shift of the halo mass range is partly driven by the evolution of the halo mass function which reduces the number density of massive halos at higher $z$. Furthermore, halos of a given mass have a higher growth rate at higher redshift (Behroozi & Silk, 2015) and thus, for a non-evolving SFE – halo mass relation, harbor brighter galaxies ($\mathcal{L}\propto{}\mathcal{S}\dot{M}_{\rm halo}$). Consequently, a larger number of galaxies in low mass halos is UV detectable at higher $z$.
@@
During cosmic reionization ($z\sim{}6-10$) the neutral hydrogen component of the intergalactic medium (IGM) is converted into its ionized state (e.g., Fan et al. 2006; Robertson et al. 2010). The primary physical driver of this phase transition is thought to be Lyman-continuum radiation escaping low-mass galaxies (Madau et al., 1999; Gnedin, 2000; Robertson et al., 2015; Mascia et al., 2023; Simmonds et al., 2023b; Atek et al., 2024; Dayal et al., 2024). In this section, we will explore the implications of the theoretical model implemented as described in Section 3.5 by comparing its predictions with available observational constraints on the ionization of the IGM. In particular, by linking the predicted UV luminosity to the observed ionized hydrogen fraction of the Universe, we will constrain the product of the escape fraction $f_{\rm esc}$ of ionizing radiation and the ionizing photon production efficiency $\xi_{\rm ion}$ of faint galaxies.
@@
Figure 11: Reionization history predicted by the theoretical model for a non-evolving star formation efficiency – halo mass relation inferred from FIREboxHR. In each panel, results are shown for a range of values for $f_{\rm esc}\xi_{\rm ion}$, see legend, and for a limiting magnitude of $M_{\rm UV}=-12$. (Top) Volume-weighted ionized hydrogen fraction predicted by the model. Overplotted are observational constraints based on the Lyman-$\alpha$ dark pixel fraction (DP; McGreer et al. 2015), gap-peak statistics in quasar absorption spectra (GP; Gallerani et al. 2007), Lyman-$\alpha$ damping wing absorption of quasars (QDW; Schroeder et al. 2013; Greig et al. 2017; Bañados et al. 2018; Greig et al. 2022), Lyman-$\alpha$ damping wing absorption of galaxies (GDW; Hsiao et al. 2023; Umeda et al. 2023), clustering of Lyman-$\alpha$ emitters (LAE-C; Ouchi et al. 2010; Sobacchi & Mesinger 2015), the fractions and equivalent width distributions of Lyman-$\alpha$ emitters (LAE-EW; Schenker et al. 2014; Mason et al. 2019; Bolan et al. 2021; Bruton et al. 2023), and Lyman-$\alpha$ opacity statistics (LAO; Gaikwad et al. 2023), see legend. Upper and lower limits are indicated by arrows and correspond to the 95 percent confidence level, if reported, otherwise to the 68 percent confidence level. Horizontal errorbars indicate the redshift range of the observational sample. Vertical errorbars show uncertainties in the estimated ionized volume fraction at the 68 percentile confidence level. The blue circle at the center of the upper panel indicates the mid-point of re-ionization at $z_{\rm re}=7.57$ as predicted by the theoretical model for $f_{\rm esc}\xi_{\rm ion}=10^{24.74}\,{\rm Hz\,erg}^{-1}$. (Bottom) Optical depth of cosmic microwave background (CMB) photons due to Thomson scattering off free electrons as predicted by the model. The values for $f_{\rm esc}\xi_{\rm ion}$ are chosen to match the measured optical depth by the Planck Collaboration (dotted line; Aghanim et al. 2020), obtained from CMB power spectra combined with CMB lensing reconstruction and baryonic acoustic oscillations, and the boundaries of the 68 and 95 percentile confidence interval (gray shaded regions). The model is in good agreement with observational data provided $f_{\rm esc}\xi_{\rm ion}$ lies in the range $\sim{}10^{24.5}-10^{24.7}\,{\rm Hz\,erg}^{-1}$.
@@
Given $f_{\rm esc}\,\xi_{\rm ion}$ and a cosmic UV luminosity density $\rho_{\rm UV}(z)$ provided by our theoretical model, we solve eqs. (17, 18) numerically with the scipy.integrate.solve_ivp function starting from a small, positive value for $Q$ at $t=0$ (we use $10^{-7}$). Values of $Q>1$ are clipped to $Q=1$. A limiting magnitude of $M_{\rm UV,lim}=-12$ is used to calculate $\rho_{\rm UV}(z)$. The reionization history predicted by this model is insensitive to the choice of $M_{\rm UV,lim}\geq{}-14$.
@@
where $\sigma_{\rm T}$ is the Thomson cross section, $c$ is the speed of light, $H(z)$ is the Hubble parameter, $Q(z)$ is the reionization history computed numerically via eqs. (17, 18), $Y=0.243$ is the cosmic Helium mass fraction (Aver et al., 2015; Peimbert et al., 2016; Aghanim et al., 2020), and $\eta=1$ ($\eta=2$) assumes singly (doubly) ionized Helium at $z\geq{}4$ ($z<4$). We then find a value for $f_{\rm esc}\,\xi_{\rm ion}$ such that $\tau_{\rm e}(z=14)=\tau_{\rm e,Planck}$ with the help of the scipy.optimize.root_scalar function. To represent the uncertainty in the Planck 2018 measurement, we also constrain $f_{\rm esc}\,\xi_{\rm ion}$ such that $\tau_{\rm e}(z=14)$ equals $\tau_{\rm e,Planck}\pm{}0.0071$ and $\tau_{\rm e,Planck}\pm{}2\times{}0.0071$.
Fig. 11 compares the reionization history $Q(z)$ predicted by our theoretical model with a variety of observational data based on Lyman-$\alpha$ dark pixel fractions (McGreer et al. 2015), gap-peak statistics in quasar absorption spectra (Gallerani et al., 2007), Lyman-$\alpha$ damping wing absorption of quasars and galaxies (Schroeder et al., 2013; Greig et al., 2017; Bañados et al., 2018; Greig et al., 2022; Hsiao et al., 2023; Umeda et al., 2023), clustering of Lyman-$\alpha$ emitters (Ouchi et al., 2010; Sobacchi & Mesinger, 2015), fractions and equivalent width distributions of Lyman-$\alpha$ emitters (Schenker et al., 2014; Mason et al., 2019; Bolan et al., 2021; Bruton et al., 2023), and Lyman-$\alpha$ opacity statistics (Gaikwad et al., 2023). The reionization history predicted by the theoretical model is in good agreement with observational data for $f_{\rm esc}\,\xi_{\rm ion}$ values that result in optical depths close to $\tau_{\rm e,Planck}$. Once possible exception is the lower limit of 0.12 at $z=10.6$ reported by Bruton et al. (2023) which, however, conflicts with the upper limit of 0.1 at $z=10.17$ reported by Hsiao et al. (2023). Replacing eq. (18) with the simpler ansatz $\gamma=\delta=1$ leads to reionization histories that are near identical to those shown in the figure when $Q<0.7$ but also to a much faster evolution of the ionized hydrogen fraction during the final stages of reionization.
@@
result in good agreement with most observational data. To convert constraints on $\zeta$ into an estimate of the escape fraction, we need to divide by the ionizing photon production efficiency $\xi_{\rm ion}$. At $z\sim{}4-6$ the efficiency $\lg{}(\xi_{\rm ion}\,[{\rm Hz}\,{\rm erg}^{-1}])$ is $\sim{}25-25.5$ (e.g., Robertson et al. 2013; Bouwens et al. 2016a; Naidu et al. 2020; Castellano et al. 2023; Saldana-Lopez et al. 2023; Simmonds et al. 2023a), with a similar range in ionizing efficiency also predicted by numerical models (e.g., Wilkins et al. 2016; Ceverino et al. 2019; Seeyave et al. 2023). Recent estimates with JWST, however, suggest somewhat larger values up to $25.8$ in galaxies at $z\sim{}7-9$ likely as a result of their younger ages and lower metallicities (Tang et al., 2023; Fujimoto et al., 2023; Endsley et al., 2023b; Atek et al., 2024; Simmonds et al., 2023b). Adopting the upper value of $\lg{}(\xi_{\rm ion}\,[{\rm Hz}\,{\rm erg}^{-1}])=25.8$, we infer escape fractions of $f_{\rm esc}=0.05-0.08$, while lower ionization efficiencies of $25.4$ and $25.0$ would imply correspondingly larger escape fractions of $f_{\rm esc}=0.12-0.20$ and $f_{\rm esc}=0.30-0.50$.
@@
In summary, a non-evolving SFE – halo halo mass relation produces an ionization history in agreement with observations provided $\lg{}(f_{\rm esc}\,\xi_{\rm ion}[{\rm Hz}\,{\rm erg}^{-1}])$ is about $24.5-24.7$. Large ionizing efficiencies $\lg{}(\xi_{\rm ion}\,[{\rm Hz}\,{\rm erg}^{-1}])\sim{}25.8$ thus imply small escape fractions of $\sim{}5-8\%$, see also Muñoz et al. (2024). Vice versa, larger escape fractions of $\sim{}20\%$ as predicted by FIRE-2 zoom-in simulations at $z\sim{}6-10$ (Ma et al., 2020) require lower average ionization efficiencies of $25.2-25.4$ (cf. Endsley et al. 2023a) to avoid overproducing ionizing photons resulting in too early an reionization of the Universe. Future measurements of the ionization efficiency during $z\sim{}6-10$ for large representative samples with JWST should provide more clarity.
@@ 4 Summary and Conclusions
  • Further inspection of the theoretical model reveals that the comparably gradual evolution of the cosmic UV luminosity density is a consequence of the weak mass dependence (and the redshift independence) of the SFE – halo mass relation in the intermediate halo mass regime (Fig. 10). When reducing the slope of the relation while keeping the UV luminosity density at $z=6$ fixed, galaxies in $M_{\rm halo}\lesssim{}10^{10.7}\,M_{\odot}$ halos tend to become more UV luminous. Also, additional galaxies in low mass halos may exceed the adopted UV magnitude limit and thus contribute to the UV luminosity density. The resulting increase in the cosmic UV luminosity density is stronger at higher $z$ as the typical halo mass harboring galaxies contributing to the cosmic UV luminosity density becomes smaller. Hence, a lower (i.e., shallower) slope of the SFE – halo mass relation leads to a more gradual evolution of the UV luminosity density.
@@ Acknowledgements
RF acknowledges financial support from the Swiss National Science Foundation (grant no PP00P2_157591, PP00P2_194814, and 200021_188552). MBK acknowledges support from NSF CAREER award AST-1752913, NSF grants AST-1910346 and AST-2108962, NASA grant 80NSSC22K0827, and HST-GO-16686, HST-AR-17028, and HST-AR-17043 from the Space Telescope Science Institute, which is operated by AURA, Inc., under NASA contract NAS5-26555. DK was supported by NSF grant AST-210834. AL was supported by NASA grant 80NSSSC20K1469. GS was supported by a CIERA Postdoctoral Fellowship. JSB was supported by NASA grant 80NSSC22K0827. PO acknowledges funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. This work was supported in part by a Simons Investigator award from the Simons Foundation (EQ) and by NSF grant AST-2107872. CAFG was supported by NSF through grants AST-1715216, AST-2108230, and CAREER award AST-1652522; by NASA through grant 17-ATP17-0067; by STScI through grant HST-AR-16124.001-A; and by the Research Corporation for Science Advancement through a Cottrell Scholar Award. This research was supported in part by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project #562 (First Light at Cosmic Dawn: Exploiting the James Webb Space Telescope Revolution) and by grant NSF PHY-1748958 to the Kavli Institute for Theoretical Physics (KITP). This work also benefited from discussions during the Nordita Program "Cosmic Dawn at High Latitudes". We acknowledge PRACE for awarding us access to MareNostrum at the Barcelona Supercomputing Center (BSC), Spain. This research was partly carried out via the Frontera computing project at the Texas Advanced Computing Center. Frontera is made possible by National Science Foundation award OAC-1818253. This work was supported in part by an allocation from the Swiss National Supercomputing Centre (CSCS) under project IDs s697 and s698. We acknowledge access to Piz Daint and Alps/Eiger at the Swiss National Supercomputing Centre, Switzerland under the University of Zurich’s share with the project ID uzh18. This work made use of infrastructure services provided by S3IT (www.s3it.uzh.ch), the Service and Support for Science IT team at the University of Zurich. All plots were created with the Matplotlib library for visualization with Python (Hunter, 2007). This research has made use of NASA’s Astrophysics Data System.

### v2

@@ 3.4 A theoretical model for the UV luminosity function and density at $z>6$
In the left panel of Fig. 10, we plot ${\rm d}\rho_{\rm UV}/{\rm d}\lg{}M_{\rm halo}$ as function of mass and redshift as predicted by our theoretical model for a mass-dependent scatter. According the figure, the majority of the cosmic UV luminosity at $z\sim{}6-15$ is produced by galaxies in halos of moderate mass ($M_{\rm halo}\sim{}10^{9}-10^{11}\,M_{\odot}$). For instance, at $z=6$ half the UV luminosity arises in halos with $M_{\rm halo}\sim{}10^{10.3}-10^{11}\,M_{\odot}$ if a UV magnitude limit of $M_{\rm UV,lim}=-17$ is chosen. The mass range of halos responsible for the majority of the UV luminosity shifts towards lower masses with increasing redshift, e.g., it is $M_{\rm halo}\sim{}10^{9.1}-10^{9.6}\,M_{\odot}$ at $z=15$. For $M_{\rm UV,lim}=-14$, the corresponding halo mass ranges are $M_{\rm halo}\sim{}10^{9.7}-10^{10.7}\,M_{\odot}$ and $M_{\rm halo}\sim{}10^{8.8}-10^{9.3}\,M_{\odot}$. The downward shift of the halo mass range is partly driven by the evolution of the halo mass function which reduces the number density of massive halos at higher $z$. Furthermore, halos of a given mass have a higher growth rate at higher redshift (Behroozi & Silk, 2015; Rodríguez-Puebla et al., 2016) and thus, for a non-evolving SFE – halo mass relation, harbor brighter galaxies ($\mathcal{L}\propto{}\mathcal{S}\dot{M}_{\rm halo}$). Consequently, a larger number of galaxies in low mass halos is UV detectable at higher $z$.
@@
During cosmic reionization ($z\sim{}6-10$), the neutral hydrogen component of the intergalactic medium (IGM) is converted into its ionized state (e.g., Fan et al. 2006; Robertson et al. 2010). The primary physical driver of this phase transition is thought to be Lyman-continuum radiation escaping low-mass galaxies (Madau et al., 1999; Gnedin, 2000; Robertson et al., 2015; Mascia et al., 2023; Simmonds et al., 2023b; Atek et al., 2024; Dayal et al., 2024). In this section, we will explore the implications of the theoretical model implemented as described in Section 3.5 by comparing its predictions with available observational constraints on the ionization of the IGM. Specifically, by linking the predicted UV luminosity to the observed ionized hydrogen fraction of the Universe, we will constrain the product of the escape fraction $f_{\rm esc}$ of ionizing radiation and the ionizing photon production efficiency $\xi_{\rm ion}$ of faint galaxies.
@@
Figure 11: Reionization history predicted by the theoretical model for a non-evolving star formation efficiency – halo mass relation inferred from FIREboxHR. In each panel, results are shown for a range of values for $f_{\rm esc}\xi_{\rm ion}$, see legend, and for a limiting magnitude of $M_{\rm UV}=-12$. (Top) Volume-weighted ionized hydrogen fraction predicted by the model. Overplotted are observational constraints based on the Lyman-$\alpha$ dark pixel fraction (DP; McGreer et al. 2015), gap-peak statistics in quasar absorption spectra (GP; Gallerani et al. 2007), Lyman-$\alpha$ damping wing absorption of quasars (QDW; Schroeder et al. 2013; Greig et al. 2017; Bañados et al. 2018; Greig et al. 2022), Lyman-$\alpha$ damping wing absorption of galaxies (GDW; Hsiao et al. 2024; Umeda et al. 2024), clustering of Lyman-$\alpha$ emitters (LAE-C; Ouchi et al. 2010; Sobacchi & Mesinger 2015), the fractions and equivalent width distributions of Lyman-$\alpha$ emitters (LAE-EW; Schenker et al. 2014; Mason et al. 2019; Bolan et al. 2021; Bruton et al. 2023), and Lyman-$\alpha$ opacity statistics (LAO; Gaikwad et al. 2023), see legend. Upper and lower limits are indicated by arrows and correspond to the 95 percent confidence level, if reported, otherwise to the 68 percent confidence level. Horizontal errorbars indicate the redshift range of the observational sample. Vertical errorbars show uncertainties in the estimated ionized volume fraction at the 68 percentile confidence level. The blue circle at the center of the upper panel indicates the mid-point of re-ionization at $z_{\rm re}=7.57$ as predicted by the theoretical model for $f_{\rm esc}\xi_{\rm ion}=10^{24.74}\,{\rm Hz\,erg}^{-1}$. (Bottom) Optical depth of cosmic microwave background (CMB) photons due to Thomson scattering off free electrons as predicted by the model. The values for $f_{\rm esc}\xi_{\rm ion}$ are chosen to match the measured optical depth by the Planck Collaboration (dotted line; Aghanim et al. 2020), obtained from CMB power spectra combined with CMB lensing reconstruction and baryonic acoustic oscillations, and the boundaries of the 68 and 95 percentile confidence interval (gray shaded regions). The model is in good agreement with observational data provided $f_{\rm esc}\xi_{\rm ion}$ lies in the range $\sim{}10^{24.5}-10^{24.7}\,{\rm Hz\,erg}^{-1}$.
@@
Given $f_{\rm esc}\,\xi_{\rm ion}$ and a cosmic UV luminosity density $\rho_{\rm UV}(z)$ provided by our theoretical model, we solve eqs. (19, 20) numerically with the scipy.integrate.solve_ivp function starting from a small, positive value for $Q$ at $t=0$ (we use $10^{-7}$). Values of $Q>1$ are clipped to $Q=1$. A limiting magnitude of $M_{\rm UV,lim}=-12$ is used to calculate $\rho_{\rm UV}(z)$. The reionization history predicted by this model is insensitive to the choice of $M_{\rm UV,lim}\geq{}-14$.
@@
where $\sigma_{\rm T}$ is the Thomson cross section, $c$ is the speed of light, $H(z)$ is the Hubble parameter, $Q(z)$ is the reionization history computed numerically via eqs. (19, 20), $Y=0.243$ is the cosmic Helium mass fraction (Aver et al., 2015; Peimbert et al., 2016; Aghanim et al., 2020), and $\eta=1$ ($\eta=2$) assumes singly (doubly) ionized Helium at $z\geq{}4$ ($z<4$). We then find a value for $f_{\rm esc}\,\xi_{\rm ion}$ such that $\tau_{\rm e}(z=14)=\tau_{\rm e,Planck}$ with the help of the scipy.optimize.root_scalar function. To represent the uncertainty in the Planck 2018 measurement, we also constrain $f_{\rm esc}\,\xi_{\rm ion}$ such that $\tau_{\rm e}(z=14)$ equals $\tau_{\rm e,Planck}\pm{}0.0071$ and $\tau_{\rm e,Planck}\pm{}2\times{}0.0071$.
Fig. 11 compares the reionization history $Q(z)$ predicted by our theoretical model with a variety of observational data based on Lyman-$\alpha$ dark pixel fractions (McGreer et al. 2015), gap-peak statistics in quasar absorption spectra (Gallerani et al., 2007), Lyman-$\alpha$ damping wing absorption of quasars and galaxies (Schroeder et al., 2013; Greig et al., 2017; Bañados et al., 2018; Greig et al., 2022; Hsiao et al., 2024; Umeda et al., 2024), clustering of Lyman-$\alpha$ emitters (Ouchi et al., 2010; Sobacchi & Mesinger, 2015), fractions and equivalent width distributions of Lyman-$\alpha$ emitters (Schenker et al., 2014; Mason et al., 2019; Bolan et al., 2021; Bruton et al., 2023), and Lyman-$\alpha$ opacity statistics (Gaikwad et al., 2023). The reionization history predicted by the theoretical model is in good agreement with observational data for $f_{\rm esc}\,\xi_{\rm ion}$ values that result in optical depths close to $\tau_{\rm e,Planck}$. Once possible exception is the lower limit of 0.12 at $z=10.6$ reported by Bruton et al. (2023) which, however, conflicts with the upper limit of 0.1 at $z=10.17$ reported by Hsiao et al. (2024). Replacing Eq. (20) with the simpler ansatz $\gamma=\delta=1$ leads to reionization histories that are near identical to those shown in the figure when $Q<0.7$ but also to a much faster evolution of the ionized hydrogen fraction during the final stages of reionization.
@@
result in good agreement with most observational data. To convert constraints on $\zeta$ into an estimate of the escape fraction, we need to divide by the ionizing photon production efficiency $\xi_{\rm ion}$. At $z\sim{}4-6$ the efficiency $\lg{}(\xi_{\rm ion}\,[{\rm Hz}\,{\rm erg}^{-1}])$ is $\sim{}25-25.5$ (e.g., Robertson et al. 2013; Bouwens et al. 2016a; Naidu et al. 2020; Castellano et al. 2023; Saldana-Lopez et al. 2023; Simmonds et al. 2023a), with a similar range in ionizing efficiency also predicted by numerical models (e.g., Wilkins et al. 2016; Ceverino et al. 2019; Seeyave et al. 2023). Recent estimates with JWST, however, suggest somewhat larger values up to $25.8$ in galaxies at $z\sim{}7-9$ likely as a result of their younger ages and lower metallicities (Tang et al., 2023; Fujimoto et al., 2023; Endsley et al., 2023; Atek et al., 2024; Simmonds et al., 2023b). Adopting the upper value of $\lg{}(\xi_{\rm ion}\,[{\rm Hz}\,{\rm erg}^{-1}])=25.8$, we infer escape fractions of $f_{\rm esc}=0.05-0.08$, while lower ionization efficiencies of $25.4$ and $25.0$ would imply correspondingly larger escape fractions of $f_{\rm esc}=0.12-0.20$ and $f_{\rm esc}=0.30-0.50$.
@@
In summary, a non-evolving SFE – halo halo mass relation produces an ionization history in agreement with observations provided $\lg{}(f_{\rm esc}\,\xi_{\rm ion}[{\rm Hz}\,{\rm erg}^{-1}])$ is about $24.5-24.7$. Large ionizing efficiencies $\lg{}(\xi_{\rm ion}\,[{\rm Hz}\,{\rm erg}^{-1}])\sim{}25.8$ thus imply small escape fractions of $\sim{}5-8\%$, see also Muñoz et al. (2024). Vice versa, larger escape fractions of $\sim{}20\%$ as predicted by FIRE-2 zoom-in simulations at $z\sim{}6-10$ (Ma et al., 2020) require lower average ionization efficiencies of $25.2-25.4$ (cf. Endsley et al. 2024) to avoid overproducing ionizing photons resulting in too early an reionization of the Universe. Future measurements of the ionization efficiency during $z\sim{}6-10$ for large representative samples with JWST should provide more clarity.
@@ 4 Summary and Conclusions
  • Further inspection of the theoretical model reveals that the comparably gradual evolution of the cosmic UV luminosity density is a consequence of the weak mass dependence (and the redshift independence) of the SFE – halo mass relation in the intermediate halo mass regime (Fig. 10). When reducing the slope of the relation while keeping the UV luminosity density at $z=6$ fixed, galaxies in $M_{\rm halo}\lesssim{}10^{10.7}\,M_{\odot}$ halos tend to become more UV luminous. Also, additional galaxies in low mass halos may exceed the adopted UV magnitude limit and thus contribute to the UV luminosity density. The resulting increase in the cosmic UV luminosity density is stronger at higher $z$ as the typical halo mass harboring galaxies contributing to the cosmic UV luminosity density becomes smaller. Hence, a lower (i.e., shallower) slope of the SFE – halo mass relation in the intermediate halo mass range leads to a more gradual evolution of the UV luminosity density. In contrast, the large slopes ($\sim{}1-1.2$) adopted by previous non-evolving SFE models (Tacchella et al., 2018; Harikane et al., 2022) lead to a steep evolution of $\rho_{\rm UV}(z)$ at $z\sim{}6-14$.
@@ Acknowledgements
The authors are grateful to the referee for their thoughtful suggestions, which improved the manuscript. RF thanks Andrey Kravtsov for insightful discussions which inspired the addition of Appendix D. RF acknowledges financial support from the Swiss National Science Foundation (grants PP00P2_194814 and 200021_188552). MBK acknowledges support from NSF CAREER award AST-1752913, NSF grants AST-1910346 and AST-2108962, and HST-GO-16686, HST-AR-17028, and HST-AR-17043 from the Space Telescope Science Institute, which is operated by AURA, Inc., under NASA contract NAS5-26555. MBK and JSB acknowledge supported from NASA grant 80NSSC22K0827. CAFG was supported by NSF through grants AST-1715216, AST-2108230, and CAREER award AST-1652522; by NASA through grant 17-ATP17-0067; by STScI through grant HST-AR-16124.001-A; and by the Research Corporation for Science Advancement through a Cottrell Scholar Award. DK was supported by NSF grant AST-210834. AL was supported by NASA grant 80NSSSC20K1469. PO acknowledges funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. This work was supported in part by a Simons Investigator award from the Simons Foundation (EQ) and by NSF grant AST-2107872. GS was supported by a CIERA Postdoctoral Fellowship. This research was supported in part by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project #562 (First Light at Cosmic Dawn: Exploiting the James Webb Space Telescope Revolution) and by grant NSF PHY-1748958 to the Kavli Institute for Theoretical Physics (KITP). This work also benefited from discussions during the Nordita Program "Cosmic Dawn at High Latitudes". We acknowledge PRACE for awarding us access to MareNostrum at the Barcelona Supercomputing Center (BSC), Spain. This research was partly carried out via the Frontera computing project at the Texas Advanced Computing Center. Frontera is made possible by National Science Foundation award OAC-1818253. This work was supported in part by an allocation from the Swiss National Supercomputing Centre (CSCS) under project IDs s697 and s698. We acknowledge access to Piz Daint and Alps/Eiger at the Swiss National Supercomputing Centre, Switzerland under the University of Zurich’s share with the project ID uzh18. This work made use of infrastructure services provided by S3IT (www.s3it.uzh.ch), the Service and Support for Science IT team at the University of Zurich. All plots were created with the Matplotlib library for visualization with Python (Hunter, 2007). This research has made use of NASA’s Astrophysics Data System.
Data availability
The data supporting the plots within this article are available on reasonable request to the corresponding author. A public version of the GIZMO code is available at http://www.tapir.caltech.edu/~phopkins/Site/GIZMO.html. FIRE data releases are publicly available at http://flathub.flatironinstitute.org/fire.